# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 5.3. Uniqueness of invariant measures
The existence and uniqueness of solutions in $C([0,T];H)$ follows under Hypotheses (A)(B)(C)(D) by [56, Theorem 5.1.3].
@@
Hence, by a standard comparison principle, we have for $\alpha>2$ that
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (G) hold true. For any $T>0$, for any $\varepsilon>0$, and $K\subset H$ bounded, such that we have
$\mathbb{P}\left(\|X_{T}^{x}-u_{T}^{x}\|_{H}^{2}<\varepsilon\right)>0,$
@@
Let $T>0$, $\varepsilon>0$, $x\in K$. Note that $Y_{t}^{x}:=X_{t}^{x}-L_{t}$ is absolutely continuous and solves the non-autonomous random PDE
(5.23) $\frac{\mathrm{d}}{\mathrm{d}t}Y^{x}_{t}=A(Y^{x}_{t}+L_{t}),\quad Y_{0}=x,$
where $(L_{t})_{t\in[0,T]}$ is as in (3.4). By Lemma C.1, for any $\delta>0$, there exists an event $\Omega_{\delta}\in{\mathcal{F}}$ with positive probability such that
$\sup_{t\in[0,T]}\|L_{t}(\omega)\|_{V}<\frac{\delta}{2}$
@@
(5.24) $\begin{split}\|Y_{t}^{x}(\omega)\|_{H}^{2}=&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(Y_{s}^{x}(\omega)+L_{s}(\omega)),Y_{s}^{x}(\omega)\rangle\,\mathrm{d}s\\ =&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),X_{s}^{x}(\omega)-L_{s}(\omega)\rangle\,\mathrm{d}s\\ \leq&\|x\|^{2}_{H}-\delta_{4}\int_{0}^{t}\|A(X_{s}^{x}(\omega))\|_{V^{\ast}}\,\mathrm{d}s-2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),L_{s}(\omega)\rangle\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|^{2}_{H}-\left(\delta_{4}-2\sup_{s\in[0,t]}\|L_{s}(\omega)\|_{V}\right)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|_{H}^{2}-(\delta_{4}-\delta)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t,\quad t\in[0,T].\end{split}$
@@
for $0<\delta<\frac{\delta_{4}}{2}$. By the chain rule, Hypothesis (C),
$\|Y^{x}_{t}(\omega)-u_{t}^{x}\|^{2}_{H}=2\int_{0}^{t}\langle A(X^{x}_{s}(\omega))-A(u_{s}^{x}),X^{x}_{s}(\omega)-L_{s}(\omega)-u_{s}^{x}\rangle\,\mathrm{d}s$
and hence, for $\eta:=\frac{\delta}{2}\widetilde{C}(T)(1+\sup_{y\in K}\|y\|^{2}_{H})$,
$\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|^{2}_{H}$ $\leq\|Y^{x}_{s}(\omega)-u_{s}^{x}\|^{2}_{H}$
$\quad+\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|_{H}^{2}\int_{s}^{T}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r$
$\quad+\frac{\delta}{2}\int_{s}^{T}\|A(X^{x}_{r}(\omega))-A(u^{x}_{r})\|_{V^{\ast}}\,\mathrm{d}r$
$\leq\left(\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|_{H}^{2}+\eta\right)\int_{s}^{T}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r+\eta.$
Let $\mathbb{E}_{\delta}:=\mathbb{E}_{\mathbb{P}(\cdot\,|\,\Omega_{\delta})}$ be the expected value with respect to the conditional probability $\mathbb{P}(\cdot\,|\,\Omega_{\delta})$. Let $0\leq\lambda\leq\Lambda\leq T$ be stopping times. We get that
$\mathbb{E}_{\delta}\left[\sup_{\lambda\leq\tau\leq\Lambda}\|Y^{x}_{\tau}-u_{\tau}^{x}\|^{2}_{H}\right]$
$\leq$ $\left(\mathbb{E}_{\delta}\left[\|Y^{x}_{\lambda}-u_{\lambda}^{x}\|^{2}_{H}\right]+\eta\right)+\mathbb{E}_{\delta}\left[\left(\sup_{\lambda\leq\tau\leq\Lambda}\|Y^{x}_{\tau}-u_{\tau}^{x}\|_{H}^{2}+\eta\right)\int_{\lambda}^{\Lambda}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\right].$
@@
By the stochastic Gronwall Lemma [3, Lemma A.1],
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\geq\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{16}{\varepsilon}e^{4R}\eta+\mathbb{P}\left(\left\{C_{2}\int_{0}^{T}\|u_{t}^{x}\|_{V}^{\alpha}\|u_{t}^{x}\|_{H}^{\beta}\,\mathrm{d}t\geq R\right\}\,\bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{16}{\varepsilon}e^{4R}\eta+\frac{1}{R}\mathbb{E}_{\delta}\left[1_{\Omega_{\delta}}C_{2}\int_{0}^{T}\|u_{t}^{x}\|_{V}^{\alpha}\|u_{t}^{x}\|_{H}^{\beta}\,\mathrm{d}t\right]$
$\leq$ $\frac{8}{\varepsilon}e^{4R}\delta\widetilde{C}(T)\left(1+\sup_{y\in K}\|y\|^{2}_{H}\right)+\frac{1}{R}\frac{2C_{2}}{\delta_{1}(\beta+2)}\sup_{y\in K}\|y\|^{\beta+2}_{H}.$
@@
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\geq\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)<\frac{1}{2}.$
@@
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)\geq\frac{1}{2},$
@@
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|X^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\varepsilon\right\}\right)$
$\geq$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\left\{\sup_{0\leq t\leq T}\|L_{t}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\right)$
$=$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\Omega_{\sqrt{\varepsilon}}\right)$
$\geq$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\Omega_{\delta}\right)\geq\frac{\mathbb{P}(\Omega_{\delta})}{2}>0,$
@@
Thus we have proved the existence of a unique invariant measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ and thus the first part of Theorem 2.1 by an application of Theorem 3.10. The concentration property can be seen as follows:
@@
Assume that Hypotheses (A)(B)(C)(D) and (F) are valid. Then any invariant probability measure $\mu_{\ast}$ has finite $H$-moments of order $\alpha+\beta$. In addition,
(5.26) $\int_{H}\|z\|^{\alpha+\beta}_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}.$
@@
(5.27) $\int_{H}\|z\|_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\right)^{1/(\alpha+\beta)}.$
@@
(5.28) $\int_{H}\|z\|^{\alpha}_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c},$
@@
(5.29) $\int_{H}\|z\|_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c}\right)^{1/\alpha}.$
Here, $\gamma\in(0,\delta_{2})$ and $\mathfrak{c}=\sum_{i=1}^{5}c_{i}\geq 0$ are as in Definition 5.1.
@@
(5.30) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{H}^{\alpha+\beta}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}t\right)\wedge\ell+\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\|z\|_{H}^{\beta+2}\right)\wedge\ell.\end{split}$
@@
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\|z\|^{\beta+2}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\left(\|z\|^{\alpha+\beta}_{H}+1\right)\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z),$
and thus for $t\geq\frac{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}{2C_{2}}$,
$\left(1-\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\wedge\ell.$
For $\alpha=2$, we get for $t>0$ that
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\delta_{1}(\beta+2)}\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z),$
and thus $t\geq\frac{c_{0}^{\alpha}\delta_{1}(\beta+2)}{2C_{2}}$,
$\left(1-\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\delta_{1}(\beta+2)}\right)\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell.$
For $\alpha\geq 2$, and for any $\varepsilon>0$, we can find a large enough $t>0$, such that
(5.33) $\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\varepsilon.$
@@
(5.34) $\int_{H}\|z\|^{\alpha+\beta}_{H}\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}+\varepsilon.$
@@
For $\beta=0$, (5.18) implies for $\ell\in\mathbb{N}$, $z\in H$, $t\geq 0$,
(5.35) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{V}^{\alpha}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c}t\right)\wedge\ell+\left(\frac{C_{2}}{\lambda_{0}\delta_{1}}\|z\|_{H}^{\beta+2}\right)\wedge\ell.\end{split}$

### v4

@@ 5.3. Uniqueness of invariant measures
The existence and uniqueness of solutions in $C([0,T];H)$ follows under Hypotheses (A)(B)(C)(D) by [70, Theorem 5.1.3].
@@
Hence, by a standard comparison principle, see e.g. [96], we have for $\alpha>2$ that
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F)(G), and (H) hold true. Assume that in Hypothesis (C) $\rho\equiv 0$ or $\eta\equiv 0$ (locally monotone case). For any $T>0$, for any $\varepsilon>0$, and $K\subset H$ bounded, such that we have
$\mathbb{P}\left(\|X_{T}^{x}-u_{T}^{x}\|_{H}^{2}\leq\varepsilon\right)>0,$
@@
W.l.o.g. $\eta\equiv 0$. Let $T>0$, $\varepsilon>0$, $x\in K$. Note that $Y_{t}^{x}:=X_{t}^{x}-B\,W_{t}$, $t\in[0,T]$, is absolutely continuous and solves the non-autonomous random PDE
(5.20) $\frac{\mathrm{d}}{\mathrm{d}t}Y^{x}_{t}=A(Y^{x}_{t}+B\,W_{t}),\quad Y_{0}=x.$
By Hypothesis (G), for any $\delta>0$, there exists an event $\Omega_{\delta}\in{\mathcal{F}}$ with positive probability such that
$\sup_{t\in[0,T]}\|B\,W_{t}(\omega)\|_{V}\leq\frac{\delta}{2}$
@@
(5.21) $\begin{split}\|Y_{t}^{x}(\omega)\|_{H}^{2}=&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(Y_{s}^{x}(\omega)+B\,W_{s}(\omega)),Y_{s}^{x}(\omega)\rangle\,\mathrm{d}s\\ =&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),X_{s}^{x}(\omega)-B\,W_{s}(\omega)\rangle\,\mathrm{d}s\\ \leq&\|x\|^{2}_{H}-\delta_{4}\int_{0}^{t}\|A(X_{s}^{x}(\omega))\|_{V^{\ast}}\,\mathrm{d}s-2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),B\,W_{s}(\omega)\rangle\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|^{2}_{H}-\left(\delta_{4}-2\sup_{s\in[0,t]}\|B\,W_{s}(\omega)\|_{V}\right)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|_{H}^{2}-(\delta_{4}-\delta)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t,\quad t\in[0,T].\end{split}$
@@
for $0<\delta<\frac{\delta_{4}}{2}$. By the chain rule, Hypothesis (C), and a similar argument as in the proof of Proposition 5.3,
$\|Y^{x}_{t}(\omega)-u_{t}^{x}\|^{2}_{H}$
$=$ $2\int_{0}^{t}\langle A(X^{x}_{s}(\omega))-A(u_{s}^{x}),X^{x}_{s}(\omega)-B\,W_{s}(\omega)-u_{s}^{x}\rangle\,\mathrm{d}s$
$\leq$ $\int_{0}^{t}\left(2\|Y^{x}_{s}(\omega)-u_{s}^{x}\|_{H}^{2}+2\|B\,W_{s}(\omega)\|_{H}^{2}\right)\left(-\delta_{2}+C_{2}\|u_{s}^{x}\|_{V}^{\alpha}\|u_{s}^{x}\|_{H}^{\beta}\right)\,\mathrm{d}s$
$+\|A(X^{x}(\omega))-A(u^{x})\|_{L^{1}([0,T];V^{\ast})}\|L(\omega)\|_{L^{\infty}([0,T];V)}$
$\leq$ $\int_{0}^{t}\left(2\|Y^{x}_{s}(\omega)-u_{s}^{x}\|_{H}^{2}+\frac{\delta^{2}}{2}\right)C_{2}\|u_{s}^{x}\|_{V}^{\alpha}\|u_{s}^{x}\|_{H}^{\beta}\,\mathrm{d}s$
$+\frac{\delta}{2}\widetilde{C}(T)\left(1+\sup_{y\in K}\|y\|^{2}_{H}\right).$
@@
Let $\mathbb{E}_{\delta}:=\mathbb{E}_{\mathbb{P}(\cdot\,|\,\Omega_{\delta})}$ be the expected value with respect to the conditional probability $\mathbb{P}(\cdot\,|\,\Omega_{\delta})$. Set $\eta:=\frac{1}{2}\widetilde{C}(T)\left(1+\sup_{y\in K}\|y\|^{2}_{H}\right)$. Now, by the stochastic Gronwall inequality [40, Theorem 1.2], and by the Markov inequality, and by Corollary 5.2,
$\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}>\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{4}{\varepsilon}e^{R}\mathbb{E}\left[\frac{\delta^{2}}{2}C_{2}\int_{0}^{T}\|u_{s}^{x}\|_{V}^{\alpha}\|u_{s}^{x}\|_{H}^{\beta}\,\mathrm{d}s+\delta\eta\right]+\mathbb{P}\left(\left\{C_{2}\int_{0}^{T}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}t>R\right\}\,\bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{4}{\varepsilon}e^{R}\left(\frac{\delta^{2}C_{2}}{\delta_{1}(\beta+2)}\sup_{y\in K}\|y\|_{H}^{\beta+2}+\delta\eta\right)+\frac{1}{R}\mathbb{E}_{\delta}\left[\mathbbm{1}_{\Omega_{\delta}}C_{2}\int_{0}^{T}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}t\right]$
$\leq$ $\frac{4}{\varepsilon}e^{R}\left(\frac{\delta^{2}C_{2}}{\delta_{1}(\beta+2)}\sup_{y\in K}\|y\|_{H}^{\beta+2}+\frac{\delta}{2}\widetilde{C}(T)\left(1+\sup_{y\in K}\|y\|^{2}_{H}\right)\right)+\frac{1}{R}\frac{2C_{2}}{\delta_{1}(\beta+2)}\sup_{y\in K}\|y\|_{H}^{\beta+2}.$
@@
$\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}>\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)<\frac{1}{2}.$
@@
$\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\leq\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)\geq\frac{1}{2},$
@@
$\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|X^{x}_{t}-u_{t}^{x}\|^{2}_{H}\leq\varepsilon\right\}\right)$
$\geq$ $\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\leq\frac{\varepsilon}{4}\right\}\cap\left\{\sup_{0\leq t\leq T}\|B\,W_{t}\|^{2}_{H}\leq\frac{\varepsilon}{4}\right\}\right)$
$=$ $\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\leq\frac{\varepsilon}{4}\right\}\cap\Omega_{\sqrt{\varepsilon}}\right)$
$\geq$ $\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\leq\frac{\varepsilon}{4}\right\}\cap\Omega_{\delta}\right)\geq\frac{\mathbb{P}(\Omega_{\delta})}{2}>0,$
@@
We have proved the existence of a unique invariant measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ and thus Theorem 2.1 by an application of Theorem 3.12. The concentration property can be seen as follows:
@@
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. Then any invariant probability measure $\mu_{\ast}$ has finite $H$-moments of order $\alpha+\beta$. In addition,
(5.23) $\int_{H}\|z\|^{\alpha+\beta}_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}.$
@@
(5.24) $\int_{H}\|z\|_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)^{1/(\alpha+\beta)}.$
@@
(5.25) $\int_{H}\|z\|^{\alpha}_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{c_{1}+c_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}},$
@@
(5.26) $\int_{H}\|z\|_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{c_{1}+c_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}}\right)^{1/\alpha}.$
Here, the constants are as in Hypotheses (B)(C)(D), and (H), respectively.
@@
(5.27) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{H}^{\alpha+\beta}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}t\right)\wedge\ell+\left(\frac{2}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\|z\|_{H}^{\beta+2}\right)\wedge\ell.\end{split}$
@@
$\leq$ $\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\wedge\ell+\frac{1}{t}\frac{2}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\|z\|^{\beta+2}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\wedge\ell+\frac{1}{t}\frac{2}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\left(\|z\|^{\alpha+\beta}_{H}+1\right)\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z),$
and thus for $t\geq\frac{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}{2}$,
$\left(1-\frac{1}{t}\frac{2}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\wedge\ell+\frac{1}{t}\frac{2}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}(1\wedge\ell).$
For any $\varepsilon>0$, we can find a large enough $t>0$, such that
(5.30) $\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\wedge\ell+\varepsilon.$
@@
(5.31) $\int_{H}\|z\|^{\alpha+\beta}_{H}\mu_{\ast}(\mathrm{d}z)\leq\frac{2(c_{1}+c_{2})}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}+\varepsilon.$
@@
For $\alpha=2$ and $\beta=0$, (5.17) implies for $\ell\in\mathbb{N}$, $z\in H$, $t\geq 0$,
(5.32) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{V}^{\alpha}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{c_{1}+c_{2}}{c_{0}^{2}\lambda_{0}\delta_{1}}\right)t\wedge\ell+\left(\frac{1}{c_{0}^{\alpha}\lambda_{0}\delta_{1}}\|z\|_{H}^{2}\right)\wedge\ell.\end{split}$